(x^2+8)/(x^2-25)
find the value of x given the expression.
1
Expert's answer
2011-12-15T07:27:03-0500
(x^2+8)/(x^2-25) =0 => x^2+8 =0 => x^2 = -8; x=sqrt{-8} = i2 sqrt{2}; where i = sqrt{-1}